OpenStudy (anonymous):

(3u^2v^5w^-4)^3/(2uv^-3w^2)^4

OpenStudy (dumbcow):

\[\frac{(3u^{2}v^{5}w^{-4})^{3}}{(2uv^{-3}w^{2})^{4}} = \frac{27u^{6}v^{15}w^{-12}}{16u^{4}v^{-12}w^{8}} = \frac{27u^{2}v^{27}}{16w^{20}}\]